Dana is in charge of planning for her sorority's social events. She is currently determining the best way to supply the posters needed for upcoming social events. In the upcoming winter semester, she expects to need 400 posters. Following that, in the spring, she will need 600 posters. Finally, the summer semester will require 250 posters. If she makes a poster ahead of time, she can store it in a locker for $1.50 per poster per semester. She has also decided that the event must have the planned number of posters; Back-orders are not allowed. To meet these demands, utilize the Transportation Model and the following information: make the posters 'in-house' by using members of the social committee during their regular meeting hours. Based on time and effort, each poster would cost about $12.50, and her team could make 150 posters per semester ask the social committee to stay late to work on posters. This would cost about $17 per poster, based on the materials, time and a 'bribe' of pizza and drinks. This could generate up to 250 posters per semester. ask another club to help out. She would have to pay this club $20 per poster, and they could create up to 500 per semester What are the lowest possible costs for Dana? How many posters are made by: the social committee during regular meetings? the social committee when staying late? the other clubs? Before she finishes creating her plan, Dana finds out that she has 50 posters already completed and ready to use. What is the impact on her total costs?
Answer #1
In order to formulate the problem as a transportation model, first create the parameter table.
Rows of the parameter table show supply period and columns indicate demand period
Regular cost of production in a supply period to fulfil the demand of same period is as given in the problem. For example regular production cost in period 1 to fulfil the demand of period 1 is $ 12.5 and overtime cost in the same supply and demand period is $ 17. Subcontract cost is $ 20

Regular time Cost of fulfilling demand of n period later than the current period increases by n times the carrying cost. For example, cost of fulfilling demand of period 3 (2 period later) from overtime time supply of period 1 = 17+2*1.5 = 20

Total supply exceeds the Net demand. Therefore, there is a dummy demand node to make it a balanced transportation problem

Resulting parameter table is as follows
Parameter Table
Supply winter spring summer dummy Supply
Initial Inventory 0 1.5 3 0 0
winter Regular 12.5 14 15.5 0 150
winter Overtime 17 18.5 20 0 250
winter Other club 20 21.5 23 0 500
spring Regular 12.5 14 0 150
spring Overtime 17 18.5 0 250
spring Other club 20 21.5 0 500
summer Regular 12.5 0 150
summer Overtime 17 0 250
summer Other club 20 0 500
Net Demand 400 600 250 1,450

Create spreadsheet model as follows:

I J K L M M24 - fx A B C D 1 Demand & Supply winter spring summer 2 Demand 400 600 250 3 Capacity 4. Regular 1 150 150 150 5

EXCEL FORMULAS:

Cell Formula Copy to
M12 =SUM(I12:L12) M12:M21
I22 =SUM(I12:I21) I22:L22
M24 =SUMPRODUCT(B12:E21,I12:L21)

Enter following Solver Parameters :

S Solver Parameters Set Objective: SM$24 To: O Max Min O Value Of: O By Changing Variable Cells: $1$12:$L$15, $J$16:$L$18,SK$

Click Solve to generate the solution:

I J K L M N O P M24 @ fx =SUMPRODUCT(B12:E21,112:L21) A B C D E F G H 1 Demand & Supply winter spring summer Cost data 2 Dema

The optimal production plan is:

Production winter spring summer
Regular 150 150 150
Overtime 250 250 100
Other club 200

Lowest possible costs = $ 19,825

Number of posters made by

the social committee during regular meetings = 150+150+150 = 450

the social committee when staying late = 250+250+100 = 600

the other clubs = 200

-----------------------------------------------------------------------------------------

Enter the initial inventory in cell F12 and again Run the Solver

@ fx B C D winter spring summer 400 600 250 I J K L M S Solver Results M24 A 1 Demand & Supply 2 Demand 3 Capacity 4 Regular

The new production plan is following:

Production winter spring summer
Regular 150 150 150
Overtime 250 250 100
Other club 150

Minimum total cost = $ 18,900

We see that after accounting for the initial inventory, total costs decreased by = 19825 - 18900 = $ 925